# State preparation by shallow circuits using feed forward

### v2

@@ 1 Introduction
The main part of our paper concerns new efficient state-preparation circuits for three types of states. All states corresponding to these three types are non-stabilizer states. Efficient circuits for preparing stabilizer states have been known through measurement-based quantum computing, we discuss this result in more detail in Section 3.2. The first state is a uniform superposition over an arbitrary number of states. This circuit includes the use of an exact version of Grover search as a subroutine, that turns a probabilistic circuit, with a known constant probability of success, into a deterministic circuit. We use the circuit for preparing a uniform superposition over an arbitrary number of states as a subroutine in the next two quantum state preparation protocols.
The second state is the $W$-state, the uniform superposition over all computational basis states of Hamming-weight $1$, a natural long-ranged entangled state that displays a fundamentally nonequivalent type of entanglement from the Greenberger–Horne–Zeilinger state [WState:2000], for which $\operatorname{\mathsf{LAQCC}}$-type constant-depth circuits were previously known [PhamSvore2013, Cirac:2021]. The W-state is often used as benchmark for new quantum hardware [Haffner2005, Neeley2010, GarciaPerez:2021]. A circuit for preparing the $W$-state was given in [Cirac:2021], but this implementation requires sequentially alternating measurements followed by local unitaries, which in the $\operatorname{\mathsf{LAQCC}}$ model is not considered to be of constant depth. We improve this protocol by giving an $\operatorname{\mathsf{LAQCC}}$ implementation of the $W$-state, based on a compress-uncompress method that links the one-hot and binary encoding of integers.
@@ 2.1 Complexity classes
The class $\operatorname{\mathsf{TC}}^{k}$ consists of all decision problems solvable by circuits of $\mathcal{O}((\log n)^{k})$ depth and polynomial size, and consisting of unbounded-fan-in AND-, OR- and Thresholdt-gates. A Thresholdt-gate evaluates to one if and only if the sum of the inputs is at least $t$.
@@ 2.3 Two quantum subroutines
Then, there exists a quantum amplitude amplification algorithm that makes $\mathcal{O}(\sqrt{N/m})$ queries to $O_{g}$ and prepares the quantum state $\sum_{x\in T}\left|x\right\rangle$.
@@ 3 The $\operatorname{\mathsf{LAQCC}}$ model
This section formally defines the $\operatorname{\mathsf{LAQCC}}$ model. Next, we shows that all Clifford circuits have an efficient and equivalent $\operatorname{\mathsf{LAQCC}}$ circuit. We then give quantum gates and tools constructable within $\operatorname{\mathsf{LAQCC}}$, such as the quantum fan-out gate and weighted threshold gate, and we conclude by showing that any $\operatorname{\mathsf{LAQCC}}$-circuit has an equivalent $\operatorname{\mathsf{QNC}}^{1}$-circuit.
@@ 3.3 Useful gates and routines with an $\operatorname{\mathsf{LAQCC}}$ implementation
Prior works extensively studied the fanout gate, for instance to construct a constant-depth $\text{OR}_{n}$ function with one-sided error [HoyerSpalek:2005] and with an exact implementation [TakahashiTani_CollapseOfHierarchyConstantDepthExactQuantumCircuits_2013], both assuming the fanout gate to be a native gate. The $\text{OR}_{n}$ gate also implies two other gates, as the following table shows.
@@
$\text{OR}_{n}$ $\left|y_{1}\right\rangle\ldots\left|y_{n}\right\rangle\left|x\right\rangle\mapsto\left|y_{1}\right\rangle\ldots\left|y_{n}\right\rangle\left|\text{OR}_{n}(y)\oplus x\right\rangle$ $\mathcal{O}(n\log(n))$ [TakahashiTani_CollapseOfHierarchyConstantDepthExactQuantumCircuits_2013, Theorem 1]
@@ 3.6 Complexity results for $\operatorname{\mathsf{LAQCC}}(\mathcal{Q},\mathcal{C},d)$
Note that for any non-zero $\varepsilon$, we can restrict ourselves to finite universal gate-sets. The Solovay-Kitaev theorem [Kitaev1997, Nielsen2012] says that any multi-qubit unitary can be approximated to within precision $\delta$ by a quantum circuit with size depending on $\delta$. Therefore, with a finite universal gate-set, any $\operatorname{\mathsf{LAQCC}}^{*}$ circuit with a continuous gate-set can be approximated by an $\operatorname{\mathsf{LAQCC}}^{*}$ circuit with gates from the finite set.
@@ 4 State preparation in $\operatorname{\mathsf{LAQCC}}$
In this section we consider what quantum states we can prepare using an $\operatorname{\mathsf{LAQCC}}$ circuit beyond the stabilizer states and Clifford circuits discussed in the previous section. First, we show how to create a uniform superposition of computational basis states up to size $q$, where $q$ is not a power of $2$. We then use this procedure to create $W$-states, the uniform superposition over all $n$-bitstrings of Hamming-weight $1$, and their generalization Dicke-$(n,k)$ states, the uniform superposition over all $n$-bitstrings of Hamming-weight $k$.
@@ 4.1 Uniform superposition of size $q$
A simple Hadamard gate applied to $n$ qubits prepares the uniform superposition $\frac{1}{\sqrt{2^{n}}}\sum_{i=0}^{2^{n}-1}\left|i\right\rangle$. Preparing the state $\frac{1}{\sqrt{q}}\sum_{i=0}^{q-1}\left|i\right\rangle$, the superposition up to size $q$, is already harder for arbitrary $q$.
@@ Acknowledgements
We want to thank Jonas Helsen, Joris Kattemölle, Ido Niesen, Kareljan Schoutens, Florian Speelman, Dyon van Vreumingen and Jordi Weggemans for insight full discussions. Furthermore, we would like to thank Georgios Styliaris for first mentioning how to parallelize Clifford ladder circuits. HB and MF were supported by the Dutch Ministry of Economic Affairs and Climate Policy (EZK), as part of the Quantum Delta NL programme.
@@ Appendix B Gate implementations
B.1 Equality-gate

### v5

@@ 1 Introduction
All Clifford circuits have a constant-depth $\operatorname{\mathsf{LAQCC}}$ implementation, implying that any stabilizer state can be implemented by a constant-depth $\operatorname{\mathsf{LAQCC}}$ circuit, see Section 3.2 for a proof of this statement. Efficient circuits for stabilizer states have been known already through measurement-based quantum computing. Therefore this paper focuses on the preparation of non-stabilizer states, and as a surprising result we find novel constant-depth protocols for four very natural classes of non-stabilizer states. Despite the extensive research into these four classes of non-stabilizer states and the many applications of them, no efficient constant- or low-depth state preparation protocols are known yet. We specifically consider these four classes as they are all often used as initial states in other algorithms.
The first state is a uniform superposition over an arbitrary number of states. This state finds applications in many quantum algorithms, as they often start with a uniform superposition over multiple states. This superposition is often achieved by applying Hadamard gates to every qubit due to its simplicity to prepare. Yet, the analysis of many algorithms, such as Shor’s algorithm [Shor:1997], would benefit from a different initial superposition. The circuit to prepare the uniform superposition over an arbitrary number of states uses an exact version of Grover search as a subroutine, that turns a probabilistic circuit, with a known constant probability of success, into a deterministic circuit. We use the circuit for preparing a uniform superposition over an arbitrary number of states as a subroutine in the next two quantum state preparation protocols.
The second state is the $W$-state, the uniform superposition over all computational basis states of Hamming-weight $1$, a natural long-ranged entangled state that displays a fundamentally nonequivalent type of entanglement from the Greenberger–Horne–Zeilinger state [WState:2000], for which $\operatorname{\mathsf{LAQCC}}$-type constant-depth circuits were previously known [PhamSvore2013, Cirac:2021]. The $W$-state is often used as benchmark for new quantum hardware [Haffner2005, Neeley2010, GarciaPerez:2021]. A novel way to prepare the $W$-state therefore gives a new way to benchmark different quantum devices with each other. A circuit for preparing the $W$-state was given in [Cirac:2021], but this implementation requires sequentially alternating measurements followed by local unitaries, which in the $\operatorname{\mathsf{LAQCC}}$ model is not considered to be of constant depth. We improve this protocol by giving an $\operatorname{\mathsf{LAQCC}}$ implementation of the $W$-state, based on a compress-uncompress method that links the one-hot and binary encoding of integers.
@@
We continue with a log-depth state preparation protocol for the Dicke-state for arbitrary $k$. This protocol implements an efficient transformation between the factoradic number representation and the combinatorial number representation of a positive integer. The combinatorial number representation relates directly to the Dicke state. The provided efficient transformation between number representation systems might be of independent interest.
We conclude by modifying our protocol for preparing a Dicke-state to a protocol that prepares quantum many-body scar states in constant-depth. These states have low entanglement and longer coherence times than states with similar energy density. These characteristics make many-body scar states interesting to analyze and relevant within physics. Many-body scar states appear for instance in the AKLT model [AKLT:1987, MRBAR:2018, MRB:2018] and different spin models [SI:2019, MOBFR:2020]. Known methods for preparing these states have polynomial-depth [Gustafson:2023], whereas our circuit has constant depth.
@@
Summarized in a table, we provide the following state generation protocols:
State description Width Depth Implementation
Uniform superposition mod $q$: $\frac{1}{\sqrt{q}}\sum_{i=0}^{q-1}\left|i\right\rangle$ $\mathcal{O}(\left\lceil\log^{2}q\right\rceil)$ $\mathcal{O}(1)$ Section 4.1
$W$-state: $\frac{1}{\sqrt{n}}\sum_{i=0}^{n-1}\left|e_{i}\right\rangle$ $\mathcal{O}(n\log n)$ $\mathcal{O}(1)$ Section 4.2
Dicke-$(n,k)$, $k=\mathcal{O}(\sqrt{n})$: $\binom{n}{k}^{-1/2}\sum_{x\in\{0,1\}^{n}:|x|=k}\left|x\right\rangle$ $\mathcal{O}(n^{2}\log n)$ $\mathcal{O}(1)$ Section 4.3
Dicke-$(n,k)$: $\binom{n}{k}^{-1/2}\sum_{x\in\{0,1\}^{n}:|x|=k}\left|x\right\rangle$ $\mathcal{O}(\text{poly}(n))$ $\mathcal{O}(\log n)$ Section 4.4
QMBS: $\left|S_{k}\right\rangle=\frac{1}{k!\sqrt{\mathcal{N}(n,k)}}(Q^{\dagger})^{k}\left|\Omega\right\rangle$ $\mathcal{O}(n^{2}\log n)$ $\mathcal{O}(1)$ Section 4.5
Table 1: Summary of state preparation protocols given in this paper.
In the entry for the quantum many-body scar state $Q$ denotes the raising operator and $\mathcal{N}(n,k)=\binom{n-k-1}{k}$. Section 4.5 will provide more details on the variables and the implementation.
@@ 2.1 Complexity classes
The class $\operatorname{\mathsf{TC}}^{k}$ consists of all decision problems solvable by circuits of $\mathcal{O}((\log n)^{k})$ depth and polynomial size, and consisting of unbounded-fan-in AND-, OR- and Thresholdt-gates, where a Thresholdt-gate evaluates to one if and only if the sum of the inputs is at least $t$.
@@ 2.3 Two quantum subroutines
Then, there exists a quantum amplitude amplification algorithm that makes $\mathcal{O}(\sqrt{N/m})$ queries to $O_{g}$ and prepares the quantum state $\frac{1}{\sqrt{|T|}}\sum_{x\in T}\left|x\right\rangle$.
@@ 3 The $\operatorname{\mathsf{LAQCC}}$ model
This section formally defines the $\operatorname{\mathsf{LAQCC}}$ model. Next, we show that all Clifford circuits have an efficient and equivalent $\operatorname{\mathsf{LAQCC}}$ circuit. We then give quantum gates and tools constructable within $\operatorname{\mathsf{LAQCC}}$, such as the quantum fan-out gate and weighted threshold gate, and we conclude by showing that any $\operatorname{\mathsf{LAQCC}}$-circuit has an equivalent $\operatorname{\mathsf{QNC}}^{1}$-circuit.
@@ 3.3 Useful gates and routines with an $\operatorname{\mathsf{LAQCC}}$ implementation
Prior works extensively studied the fanout gate, for instance to construct a constant-depth $\text{OR}_{n}$ function with one-sided error [HoyerSpalek:2005] and with an exact implementation [TakahashiTani_CollapseOfHierarchyConstantDepthExactQuantumCircuits_2013, Theorem 1], both assuming the fanout gate to be a native gate. The $\text{OR}_{n}$ gate also implies two other gates, as the following table shows.
@@
$\text{OR}_{n}$ $\left|y_{1}\right\rangle\ldots\left|y_{n}\right\rangle\left|x\right\rangle\mapsto\left|y_{1}\right\rangle\ldots\left|y_{n}\right\rangle\left|\text{OR}_{n}(y)\oplus x\right\rangle$ $\mathcal{O}(n\log(n))$ Appendix B.1
@@ 3.6 Complexity results for $\operatorname{\mathsf{LAQCC}}(\mathcal{Q},\mathcal{C},d)$
Note that for any non-zero $\varepsilon$, we can restrict ourselves to finite universal gate-sets. The Solovay-Kitaev theorem [Kitaev1997, nielsen_chuang_2010] says that any multi-qubit unitary can be approximated to within precision $\delta$ by a quantum circuit with size depending on $\delta$. Therefore, with a finite universal gate-set, any $\operatorname{\mathsf{LAQCC}}^{*}$ circuit with a continuous gate-set can be approximated by an $\operatorname{\mathsf{LAQCC}}^{*}$ circuit with gates from the finite set.
@@ 4 State preparation in $\operatorname{\mathsf{LAQCC}}$
In this section we consider what quantum states we can prepare using an $\operatorname{\mathsf{LAQCC}}$ circuit beyond the stabilizer states and Clifford circuits discussed in the previous section. Specifically, as mentioned in the introduction, we consider quantum states that are widely used, in other quantum algorithms, for bench marking purposes and within physics. First, we show how to create a uniform superposition of computational basis states up to size $q$, where $q$ is not a power of $2$, a state that is often used as initial state in other algorithms (including the following other state preparation protocols presented in this work). We then use this procedure to create $W$-states, the uniform superposition over all $n$-bitstrings of Hamming-weight $1$, using a compress-uncompress method. This compress-uncompress method generalizes to preparing Dicke-$(n,k)$ states for $k=\mathcal{O}(\sqrt{n})$, uniform superpositions over all $n$-bitstrings of Hamming-weight $k=\mathcal{O}(\sqrt{n})$. Dicke states find many applications, and especially the compress-uncompress approach might prove useful for entanglement distillation protocols. Preparing general Dicke-$(n,k)$ states requires a novel method to map between two integer representation systems, the factoradic representation and the combinatorial number representation. Finally, we present a state preparation protocol for quantum many-body scar states, states often used in physics, based on the Dicke-$(n,k)$ state preparation protocol for $k=\mathcal{O}(\sqrt{n})$
@@ 4.1 Uniform superposition of size $q$
The uniform superposition is often used as initial state in other quantum algorithms. A simple Hadamard gate applied to $n$ qubits prepares the uniform superposition $\frac{1}{\sqrt{2^{n}}}\sum_{i=0}^{2^{n}-1}\left|i\right\rangle$. Preparing the state $\frac{1}{\sqrt{q}}\sum_{i=0}^{q-1}\left|i\right\rangle$, the superposition up to size $q$, is already harder for arbitrary $q$.
@@ Acknowledgements
We want to thank Jonas Helsen, Joris Kattemölle, Ido Niesen, Kareljan Schoutens, Florian Speelman, Dyon van Vreumingen and Jordi Weggemans for insight full discussions. Furthermore, we would like to thank Georgios Styliaris for first mentioning how to parallelize Clifford ladder circuits. HB and MF were supported by the Dutch Ministry of Economic Affairs and Climate Policy (EZK), as part of the Quantum Delta NL programme. BL was funded by the European Union (ERC, HOFGA, 101041696). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Also supported by FCT through the LASIGE Research Unit, ref. UIDB/00408/2020 and ref. UIDP/00408/2020. NN was supported by the quantum application project of TNO. This work was supported by the Dutch Research Council (NWO/OCW), as part of the Quantum Software Consortium programme (project number 024.003.037).
@@ Appendix B Gate implementations
B.1 OR-gate
[…]
Let $m=\left\lceil\log_{2}(n+1)\right\rceil$, then the OR-reduction implements the map
$\left|x\right\rangle\left|0\right\rangle^{\otimes m}\to\left|x\right\rangle\bigotimes_{j=1}^{m}\left|\mu^{|x|}_{\phi_{j}}\right\rangle,$
where $\varphi_{j}=\frac{2\pi}{2^{j}}$ and $\left|\mu_{\varphi}^{|x|}\right\rangle=HR_{Z}(\varphi\cdot|x|)H\left|0\right\rangle$. The OR-reduced state thus depends on the weighted Hamming weight of $x$, more precisely, every $\left|\mu^{|x|}_{\phi_{j}}\right\rangle$ depends on the entire string $x$. For every $\left|\mu^{|x|}_{\phi_{j}}\right\rangle$ this requires a copy of $\left|x\right\rangle$ which can be created by using the fanout gate. The circuit applying the rotation $R_{Z}(\varphi\cdot|x|)$ consists of sequential $R_{Z}$ gates, which are diagonal and hence can be implemented in parallel by Lemma 2.8. This results in a circuit of width $\mathcal{O}(n\log(n))$ for creating the OR-reduced state.
We have for every bitstring $x\in\{0,1\}^{n}$ that
$\text{OR}_{n}(x)=\frac{1}{2^{n-1}}\sum_{a\in\{0,1\}^{n}\setminus\{0^{n}\}}\text{PA}_{n}^{a}(x),$
[…]
We now copy the $m$ qubits of the OR-reduced state $2^{m}=n$ times, using fanout gates. For each copy we require two additional auxiliary qubits. The first will hold the result of the parity computation, for which we already have a nearest-neighbor implementation. We will entangle the second auxiliary qubit using a fanout-gate to obtain a GHZ-state in these qubits.
We now compute in parallel the parity of the subsets of the inputs. For every subset we use a single copy of the inputs and we store the result in the corresponding auxiliary qubit. We then apply a controlled-$R_{Z}$ gate from the first auxiliary qubit to the second auxiliary qubit in the GHZ state. This prepared the state (omitting other registers)
$\frac{1}{\sqrt{2}}(\left|0\right\rangle^{n-1}+(-1)^{\frac{1}{2^{n-1}}\sum_{a\in\{0,1\}^{m}\setminus\{0^{m}\}}\text{PA}_{m}^{a}(x)}\left|1\right\rangle^{\otimes n-1})=\frac{1}{\sqrt{2}}(\left|0\right\rangle^{n-1}+(-1)^{\text{OR}_{n}(x)}\left|1\right\rangle^{\otimes n-1}).$
Uncomputing this final state using a fanout-gate gives a single qubit that holds the desired answer in its phase. A single Hadamard gate applied to this qubit will then give the answer in a single qubit.
Combining all steps thus gives an implementation for the OR gate using a geometrically local nearest-neighbor circuit. For more details on the implementation as well as a proof of correctness, we refer to the original proof [TakahashiTani_CollapseOfHierarchyConstantDepthExactQuantumCircuits_2013].
B.2 Equality-gate